OpenStudy (anonymous):

Please help! 64 is 4 times the difference between Sarah's age, a, and 44. Assume Sarah is older than 44. Write an equation to determine Sarah's age (a)

OpenStudy (anonymous):

Alright im thinking.

OpenStudy (anonymous):

You would have to find what can subtract 64 more than 4 times and the answer still has to be higher than 44.

OpenStudy (anonymous):

So what is 64 * 4 first?

OpenStudy (anonymous):

256 right?

OpenStudy (anonymous):

Oh wait i did something wrong, cuz 256 - (64 * 4) would equal 0, and thats not any higher than 44, sorry i did it wrong, I think i got an idea though.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

maybe you do 64*5 so it is more than 256 so when you subtract, the answer will be higher than 44, so maybe try that.

OpenStudy (anonymous):

64*5= 320, so now try 320 - (6*4, aka 256)

OpenStudy (anonymous):

(64*4)*

OpenStudy (anonymous):

I get 64 left over.

OpenStudy (anonymous):

oh wow. sorry Im not That good at math, maybe ask someone else ill tag someone for you.

OpenStudy (anonymous):

@the_ocean_girl @19rutlehop

OpenStudy (anonymous):

Sorry i coulnt help you _Lemons_

OpenStudy (anonymous):

It's alright <3

OpenStudy (anonymous):

the_ocean_girl, you got it from here?

Theoceangirl (the_ocean_girl):

I got it :)

OpenStudy (anonymous):

Okay thanks ^-^ Cya later.

Theoceangirl (the_ocean_girl):

Bye :)

Theoceangirl (the_ocean_girl):

Okay so the question is 64 is (=) 4 times the difference between (-) Sarah's age (a) and 44

OpenStudy (anonymous):

Okay, so I'm so confused by this, am I going to end up with something like, my original attempt was something like a = 4(64/4)... But I don't think that's even close to the answer. :(

OpenStudy (anonymous):

:O ur awesome.

Theoceangirl (the_ocean_girl):

We have 64= Now it says 4 times the value of A-44 So you would put a-44 in parentheses, multiply it by 4, and set it equal to 64 Now do you know what the equation would be?

OpenStudy (anonymous):

So 64 = 4(a-44) ?

Theoceangirl (the_ocean_girl):

Good!

OpenStudy (anonymous):

YUS! How do I give you a medal or something!

Theoceangirl (the_ocean_girl):

Click the "Best Response" button And welcome to OS :)

OpenStudy (anonymous):

Thank you!!!!

Theoceangirl (the_ocean_girl):

You're welcome :) If you need anymore help let me know :)

OpenStudy (anonymous):

To find Sarah's age, would I just solve for A in that equation?

Theoceangirl (the_ocean_girl):

Yep :)